# Improving the efficiency of finite-time memory erasure with potential barrier shaping

### v4

@@ Abstract
Erasure of the binary memory, 0 or $1$, is an essential step for digital computation as it involves irreversible logic operations. In the classical case, the erasure of a bit of memory is accompanied by the evolution of a minimum amount of heat set by the Landauer bound $k_{B}T\ln 2$, which can be achieved in the asymptotic limit. However, the erasure of memory needs to be completed within a finite time for practical and effective computational processes. It is observed that the higher the speed of erasure, the greater the amount of heat released, which leads to unfavorable environmental conditions. Therefore, this is a fundamental challenge to reduce the evolved heat related to finite-time memory erasure. In the present work, we address this crucial aspect in the field of information thermodynamics, where the two memory states correspond to the two wells of a bistable potential, as in the conventional cases. However, the potential is asymmetric in terms of the width of the two wells. Moreover, the two memory states are separated by a barrier that is asymmetric in structure. We examine in detail the effect of the degree of asymmetry on the success rate of the erasure process and the work done or heat released associated with it. We find that the asymmetry in the potential barrier partitioning the two memory states plays a very significant role in improving the efficiency of the erasure process, in view of the success rate and the thermodynamic costs. We retrieve the approach towards the Landauer limit in terms of the energetics involved with the erasure mechanism under the symmetric setup.
@@ I Introduction
Digital computation has become an inevitable part of our everyday life because of its implementation in the technological devices that appear to be indispensable to our daily needs. These computational processes mainly involve logical operations [Murali_prl_2009, Takeuchi_pre_2018, Shannon_1948] on the binary memories 0 and $1$ to generate outputs. Most of these basic types of logic operations, such as OR, AND, XOR, and their negation operations, are irreversible in nature [PhysRevX_Manzano_2024]. This is because there are at least two inputs that produce one single output, i.e., inputs cannot be determined uniquely by observing the output. Information is lost during the operational procedure, or it can be interpreted as the erasure of memory. Landauer first pointed out that the erasure of a classical bit of information is accompanied by an unavoidable evolution of heat, which has a lower bound of $k_{B}T\ln 2$  [Landauer_1961, landauer_nature_1998, 5391327, bennett1982thermodynamics, Mandal2012, Barbara_pra_2000, chattopadhyay2025landauer, Landi_prl, Chaves_prx_2025, Sander_pre2025], where $k_{B}$ is the Boltzmann constant and $T$ represents the temperature. This change may be understood in terms of the decrease of entropy $k_{B}\ln 2$ for the process that is the result of the conversion of the initially available two memory states, to one at the final time. As the computational devices form a closed system with their surroundings, the overall entropy change remains zero and therefore, the reduced entropy for the devices appears as a heating effect in the environment, with the value of $k_{B}T\ln 2$ [Dillenschneider_prl_2009, berut_lutz2012nature, Klaers_prl_2019]. Now, the knowledge of thermodynamics suggests that this is the interpretation of the amount of the released heat when the process is quasi-reversible, i.e., the time for the completion of the process is almost infinite. We understand that the memory erasure mechanism has to be completed in a given amount of time for realistic and effective computations. Any finite-time process for the erasure of a classical bit of memory would result in the evolution of heat, which has a greater value as compared to the Landauer bound of $k_{B}T\ln 2$ [finite_time_landauer_prl_2022, Proesmans_PhysRevLett_2020, Proesmans_pre, Dago_Ludovic_PhysRevLett_22, Vlatko_RevModPhys.81.1, hong2016experimental, Sagawa_prr2024]. The generated heat for the erasure of bits of information accumulates depending on the extent of the computational task performed on digital devices and released in the environment. This becomes an unfavorable situation for the surroundings as well as for the devices. Therefore, it is a fundamental need to lower the value of the heat released for the finite-time memory erasure process and eventually, for the practical computations. Consequently, this topic emerges as a field of prime significance in the domain of information thermodynamics [Dynamical_Landauer_Principle_prl_2025, thermodynamics_nature_2015, Prx_Esposito_2014, prl_ito_2018, Goold_2016, prx_ciliberto_2017, MARTINI2016108, wolpert2024stochastic].
@@
Now, to predict and implement this idea of managing the thermodynamic costs associated with the erasure protocols, it is required to represent the memory by some suitable physical entity and the act of erasing has to be mimicked by an apt physical setup. In most of the previous studies, the system is considered to correspond to a Brownian particle moving in a bistable potential [berut_lutz2012nature, Dillenschneider_prl_2009, Reguera_pre_2001, Mondal_pre_2010], which is in a thermal equilibrium with a heat bath. The state of the particles in the two separate wells of the bistable potential corresponds to two distinct memory states 0 and $1$. To understand the memory erasure in small systems, an ensemble of such setups is taken into account. Due to the application of a time-linear bias, which is considered an erasing force, during the time of the protocol, one type of memory is erased out of the two. The ensemble-averaged value of the work or the heat related to this mechanism provides an idea about the thermodynamic costs associated with the erasure protocol. Some crucial studies have been done to optimize the evolved heat associated with the erasure process [Proesmans_PhysRevLett_2020, Proesmans_pre, Pre_2025_ludovic, Naoto_2018, Mitchison_prl_2020, roldan2014universal_nature, Gopalkrishnan2016, Lee_prl_2022, Saito_prl_2022, Modi_prl_2021, konopik2023fundamental, Sagawa_2025_prr]. Different pathways of consideration of the underdamped dynamics [Dago_Ludovic_PhysRevLett_22, Dago_2022, aurell2012refined], time-multiplexed potentials[Talukdar_pre_2017], final local equilibrium state [Proesmans_PhysRevLett_2020, Konopik_2020_epl, ciampini2021experimental, Buffoni_pre], nonequilibrium memory states  [Konopik_2020_epl, saira_prr_2020, Sagawa_2009, Esposito_2011], etc. are considered in this regard. However, there is still no definite answer to this important question of reducing the heat systematically, by choosing a particular form of the memory or the erasure protocol [Proesmans_PhysRevLett_2020, Proesmans_pre]. In the majority of the earlier studies, the dynamics of the Brownian particle whose state represents the memory, are governed by a symmetric bistable potential, and the two states are separated by a symmetric barrier. In the present study, we consider the underlying bistable potential to be asymmetric in nature with respect to the width of the two potential wells and the degree of asymmetry can be tuned by an asymmetry parameter [innerbichler_2020, ciampini2021experimental]. The ultimate value of the asymmetry parameter converts the potential to a symmetric one. In this framework, the memory states are separated by an asymmetric barrier top, except for the symmetric form of the potential. We perform a detailed study to investigate whether and how this applied asymmetry in the potential can improve the efficiency of the erasure process, and consequently, the performance of the computational tasks. The answer is sought in terms of the erasure rate or success rate of the process and the thermodynamic costs, i.e., the average amount of work-done, evolved heat, and free energy related to it. The effect of the general asymmetry on the memory erasure procedure has been studied before [Sagawa_2009, Gavrilov_2016]. However, the understanding of the quantitative response of the efficacy of the erasure phenomenon towards the degree of asymmetry is missing. Here, we perform a methodical study to interpret how the extent of asymmetry in the potential influences the above-mentioned observables related to erasure to tune the efficiency of the process. The current work indicates a precise path for enhancing the effectiveness of the fundamental process of memory erasure in a systematic way.
@@ II Model and dynamics
We employ an asymmetric double-well potential $V(x,t)$ following the forms used in Refs. [innerbichler_2020], where the asymmetry arises from differing well widths. The potential is defined as
@@
where $h$ is the height of the barrier and $S(x)$ is a scaling function that regulates the relative width of the left and right wells, introducing asymmetry in the potential. $g(t)$ is a dimensionless, time-dependent function that modulates the barrier height during the erasure protocol. $Af(t)$ is an external force applied directly to the particle to cause transitions during the process. Therefore, the erasure mechanism is characterized by two time-dependent functions: $g(t)$, the barrier-lowering function, and $f(t)$, which facilitates the transition of the particle across the barrier, i.e., to the desired memory state, by introducing tilt to the potential. $A$ stands for the amplitude of this tilting force. The mathematical descriptions of these two functions are presented below. The function $g(t)$, which modulates the height of the potential barrier over time, lowering and raising it within a duration $t_{f}=\tau/2$, can be expressed as,
@@
At the end of the full erasure cycle, both functions $g(t)$ and $f(t)$ return to their initial values (Fig. 2), and the potential $V(x)$ is restored to its original form, as illustrated in Fig. 3.
Next, we proceed to obtain Eq. 1 in dimensionless form to present and deal with a simplified description of the dynamics. The equation has been made dimensionless by scaling the variables and parameters involved in the dynamics with appropriate factors. The dimensionless position variable is expressed as $\tilde{x}=x/l$ where $l$ is the distance between the two minima of the potential and has a numerical value equal to $1$. Time $t$ has been scaled with the factor $t_{l}=\gamma l^{2}/k_{B}T_{R}$ where $T_{R}$ is a reference temperature. The scaling factor $t_{l}$ effectively represents twice the time required for the Brownian particle to travel the distance $l$. The dimensionless time becomes $\tilde{t}=t/t_{l}$. The amplitude $A$ of the driving force $f(t)$ has been made unitless by dividing by the factor $\gamma l/\tau$, and the dimensionless quantity has the form of $\tilde{A}=A\tau/\gamma l$. The frequency $\omega$ appearing in the barrier-lowering term is multiplied by $t_{l}$ to get the scaled frequency as $\tilde{\omega}=\omega t_{l}$.
@@
Figure 1: Potential $U(x)$ with different asymmetry parameter value $c$ for a fixed barrier height of $h=1$.
@@
Figure 4: Stochastic trajectories of 100 particles during the erasure process designed in bistable potentials with barrier height $h=1$, diffusion constant $D=0.1$, $\omega=0.01$, and barrier-lowering depth $Z=0.5$. Top row: Symmetric potential ($c=1$), showing (a) incomplete erasure at low tilt $A=0.2$, and (b) successful erasure at $A=1$ with equilibrium distribution of initial memory states. Middle row: Asymmetric potential ($c=0.2$), showing (c) incomplete erasure at $A=0.1$, and (d) successful erasure at $A=0.3$ with equilibrium distribution of initial memory states. Bottom row: Asymmetric potential ($c=0.2$), showing (e) incomplete erasure at $A=0.1$, and (f) successful erasure at $A=0.3$ with nonequilibrium distribution $(50:50)$ of initial memory states. For the cases with initial equilibrium states, the thermalization time is $t_{0}=314$. The ramp-up and ramp-down times for the tilting force are $\tau_{1}=298$ and $\tau_{2}=16$, respectively, for all cases.
@@ III Results and discussions
In the classical scenario, the Landauer bound is understood starting from an initial equal distribution to the memory states, 0 and $1$, and then ultimate conversion to one unique memory state. In terms of the physical system, it implies initiating the process with equal occupancy of the Brownian particles in both wells, and then transferring all the particles through the erasure mechanism to a given well, which represents the final target memory state. Now, in the case of the symmetric bistable potential, this initial state is also the equilibrium state of the system before the start of the erasure protocol. Therefore, we can say that the Landauer limit is realized starting with the equilibrium memory. This initial equilibrium distribution can be achieved if we allow all Brownian trajectories to get adjusted with the thermal bath for a sufficiently long time [Aurell_2011_PhysRevLett, Dillenschneider_prl_2009, Esposito_2011, Patrick_pre]. During this time, an individual Brownian particle gets settled in either of the two wells in the absence of any external force. When we consider an ensemble of such particles, the occupancy in the two wells will follow an equilibrium, more precisely, a steady-state distribution, in the long-time limit. Now, there is one important point to note. Although after this thermalization period, an equal equilibrium distribution in both the memory states is attained in the case of the symmetric potential, the thermally equilibrated distribution remains different for the asymmetric potentials. It also varies depending on the degree of asymmetry in the form of the potential. This is because after this equilibration time, the distribution follows from the steady state probability of occupancy of the two wells, which is now unequal for different wells due to their disparate widths in the asymmetric cases. Also, the thermalized distribution depends on the extent of asymmetry in the potential as the asymmetry parameter $c$ governs the relative widths of the two wells, thereby influencing the steady-state probability of their occupancy. Therefore, one of the distinct features of our present study is that when we start the erasure process with equilibrium memory, the initial distribution of the memory states is unique based on the asymmetry of the underlying potential. We perform our investigations on the erasure phenomenon with these initial conditions to understand the effect of the asymmetry on it when we start with the equilibrium memory states.
@@
We infer that it is also important to analyze the effect of the asymmetry in results for the erasure mechanism when we start from the equally distributed initial memory states. This is because consideration of that initial state sets the limit of $k_{B}T\ln 2$ for the erasure of a classical bit of memory and interprets the Landauer bound. Therefore, we also carry out a separate set of investigations starting with the same condition of the equal initial distribution to the two memory states, irrespective of the structure of the potential. In these cases, we do not allow any thermalization time before the start of the protocol. We realize that this distribution will not conform to the thermalized distribution of the systems without any bias for the asymmetric cases, and can be regarded as the nonequilibrium initial condition of the memory states for those systems. Therefore, performing the studies with this initial distribution of memory states also provides us with the scope of comparisons of the findings between the cases of initial equilibrium and nonequilibrium memory conditions.

### v5

@@ Abstract
Erasure of the binary memory, 0 or $1$, is an essential step for digital computation as it involves irreversible logic operations. In the classical case, the erasure of a bit of memory is accompanied by the evolution of a minimum amount of heat set by the Landauer bound $k_{B}T\ln 2$, which can be achieved in the asymptotic limit. However, the erasure of memory needs to be completed within a finite time for practical and effective computational processes. It is observed that the higher the speed of erasure, the greater the amount of heat released, which leads to unfavorable environmental conditions. Therefore, this is a fundamental challenge to reduce the evolved heat related to finite-time memory erasure. In the present work, we address this crucial aspect in the field of information thermodynamics. We proceed by considering the physical model framework where the two memory states correspond to the two wells of a bistable potential, as in the conventional cases. However, the potential is asymmetric in terms of the width of the two wells. Moreover, the two memory states are separated by a barrier that is asymmetric in structure. This type of asymmetry models the two binary memory states which occupy different phase-space volumes, although they are energetically equivalent, in a general setup. We examine in detail the effect of the degree of asymmetry on the success rate of the erasure process and the work done or heat released associated with it. We find that the asymmetry in the width of the potential wells and the barrier partitioning them, i.e., the two memory states, plays a very significant role in improving the efficiency of the erasure process, in view of the success rate and the thermodynamic costs. Our thorough simulation study establishes the fact that one can reach below the Landauer bound in an appropriate asymmetric setup. Importantly, it develops a quantitative understanding of the deviation from the Landauer limit as a function of the degree of asymmetry of the potential governing the erasure mechanism. Moreover, through our current work, we identify the effective free energy change for the finite-time bit erasure process as a general lower bound for the work done or evolved heat even when the departure from the Landauer limit is observed. We retrieve the approach towards the Landauer limit in terms of the energetics involved with the erasure mechanism under the symmetric setup.
@@ I Introduction
Digital computation has become an inevitable part of our everyday life because of its implementation in the technological devices that appear to be indispensable to our daily needs. These computational processes mainly involve logical operations [1, 2, 3] on the binary memories 0 and $1$ to generate outputs. Most of these basic types of logic operations, such as OR, AND, XOR, and their negation operations, are irreversible in nature [4]. This is because there are at least two inputs that produce one single output, i.e., inputs cannot be determined uniquely by observing the output. Information is lost during the operational procedure, or it can be interpreted as the erasure of memory. Landauer first pointed out that the erasure of a classical bit of information is accompanied by an unavoidable evolution of heat, which has a lower bound of $k_{B}T\ln 2$  [5, 6, 7, 8, 9, 10, 11, 12, 13, 14], where $k_{B}$ is the Boltzmann constant and $T$ represents the temperature. This change may be understood in terms of the decrease of entropy $k_{B}\ln 2$ for the process that is the result of the conversion of the initially available two memory states, to one at the final time. As the computational devices form a closed system with their surroundings, the overall entropy change remains zero and therefore, the reduced entropy for the devices appears as a heating effect in the environment, with the value of $k_{B}T\ln 2$ [15, 16, 17]. Now, the knowledge of thermodynamics suggests that this is the interpretation of the amount of the released heat when the process is quasi-reversible, i.e., the time for the completion of the process is almost infinite. We understand that the memory erasure mechanism has to be completed in a given amount of time for realistic and effective computations. Any finite-time process for the erasure of a classical bit of memory would result in the evolution of heat, which has a greater value as compared to the Landauer bound of $k_{B}T\ln 2$ [18, 19, 20, 21, 22, 23, 24]. The generated heat for the erasure of bits of information accumulates depending on the extent of the computational task performed on digital devices and released in the environment. This becomes an unfavorable situation for the surroundings as well as for the devices. Therefore, it is a fundamental need to lower the value of the heat released for the finite-time memory erasure process and eventually, for the practical computations. Consequently, this topic emerges as a field of prime significance in the domain of information thermodynamics [25, 26, 27, 28, 29, 30, 31, 32].
@@
Now, to predict and implement this idea of managing the thermodynamic costs associated with the erasure protocols, it is required to represent the memory by some suitable physical entity and the act of erasing has to be mimicked by an apt physical setup. In most of the previous studies, the system is considered to correspond to a Brownian particle moving in a bistable potential [16, 15, 33, 34], which is in a thermal equilibrium with a heat bath. The state of the particles in the two separate wells of the bistable potential corresponds to two distinct memory states 0 and $1$. To understand the memory erasure in small systems, an ensemble of such setups is taken into account. Due to the application of a time-linear bias, which is considered an erasing force, during the time of the protocol, one type of memory is erased out of the two. The ensemble-averaged value of the work or the heat related to this mechanism provides an idea about the thermodynamic costs associated with the erasure protocol. Some crucial studies have been done to optimize the evolved heat associated with the erasure process [19, 20, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44]. Different pathways of consideration of the underdamped dynamics [21, 45, 46], time-multiplexed potentials[47], final local equilibrium state [19, 48, 49, 50], nonequilibrium memory states  [48, 51, 52, 53], etc. are considered in this regard. However, there is still no definite answer to this important question of reducing the heat systematically, by choosing a particular form of the memory or the erasure protocol [19, 20]. In the majority of the earlier studies, the dynamics of the Brownian particle whose state represents the memory, are governed by a symmetric bistable potential, and the two states are separated by a symmetric barrier. In the present study, we consider the underlying bistable potential to be asymmetric in nature with respect to the width of the two potential wells and the degree of asymmetry can be tuned by an asymmetry parameter [54, 49]. The ultimate value of the asymmetry parameter converts the potential to a symmetric one. In this framework, the memory states are separated by an asymmetric barrier top, except for the symmetric form of the potential. The chosen asymmetric potential of the above-mentioned form can be considered as the general setup to model binary memory states when they occupy different volumes in phase space; if they are allowed to get thermalized with the heat bath. The asymmetry parameter enables tuning of the relative phase space volumes of the two memory states. However, the two states of the binary memory remain energetically equivalent as the energy values of the two minima of the potential, representing the two memory states are the same, under the unperturbed condition. We perform a detailed study to investigate whether and how this applied asymmetry in the potential can improve the efficiency of the erasure process, and consequently, the performance of the computational tasks. The answer is sought in terms of the success rate of the process and the thermodynamic costs, i.e., the average amount of work-done, evolved heat, and free energy related to it. The effect of the general asymmetry on the memory erasure procedure has been studied before [52, 55]. However, the understanding of the quantitative response of the efficacy of the erasure phenomenon towards the degree of asymmetry is missing. Here, we perform a methodical study to interpret how the extent of asymmetry in the potential influences the above-mentioned observables related to erasure to tune the efficiency of the process. The current work indicates a precise path for enhancing the effectiveness of the fundamental process of memory erasure in a systematic way.
@@ II Model and dynamics
We employ an asymmetric double-well potential $V(x,t)$ following the forms used in Refs. [54], where the asymmetry arises from differing well widths. The potential is defined as
@@
where $h$ is the barrier height, $S(x)$ is a scaling function that regulates the relative widths of the left and right wells, $\Theta(x)$ denotes the Heaviside function. $c$ is a positive parameter that introduces asymmetry in the potential by controlling the relative widths of the two wells. Here, $g(t)$ is a dimensionless, time-dependent function that modulates the barrier height during the erasure protocol. The time-independent form of the potential $V(x)$ (corresponding to $g(t)=1$) has been shown in Fig. 1. $Af(t)$ is an external force applied directly to the particle to cause transitions during the process. Therefore, the erasure mechanism is characterized by two time-dependent functions: $g(t)$, the barrier-lowering function, and $f(t)$, which facilitates the transition of the particle across the barrier, i.e., to the desired memory state, by introducing tilt to the potential. $A$ stands for the amplitude of this tilting force. The mathematical descriptions of these two functions are presented below. The function $g(t)$, which modulates the height of the potential barrier over time, lowering and raising it within a duration $t_{f}=\tau/2$, can be expressed as,
@@
At the end of the full erasure cycle, both functions $g(t)$ and $f(t)$ return to their initial values (Fig. 2), and the potential $V(x)$ is restored to its original form, as illustrated in Fig. 3. The considered form of the barrier-lowering and the tilting force satisfies the condition of modulation of the barrier height and producing the appropriate bias during the erasure mechanism, in a simplified setup. Their functional form also fulfils the requirement of the return of the final values of the forces to their initial values to constitute a cycle. Therefore, the above design of the external control can be regarded as the elemental drive for the erasure process and has been considered conventionally, in earlier notable studies [15, 16]. Next, we proceed to obtain Eq. 1 in dimensionless form to present and deal with a simplified description of the dynamics. The equation has been made dimensionless by scaling the variables and parameters involved in the dynamics with appropriate factors. The dimensionless position variable is expressed as $\tilde{x}=x/l$ where $l$ is the distance between the two minima of the potential and has a numerical value equal to $1$. Time $t$ has been scaled with the factor $t_{l}=\gamma l^{2}/k_{B}T_{R}$ where $T_{R}$ is a reference temperature. The scaling factor $t_{l}$ effectively represents twice the time required for the Brownian particle to travel the distance $l$. The dimensionless time becomes $\tilde{t}=t/t_{l}$. The amplitude $A$ of the driving force $f(t)$ has been made unitless by dividing by the factor $\gamma l/\tau$, and the dimensionless quantity has the form of $\tilde{A}=A\tau/\gamma l$. The frequency $\omega$ appearing in the barrier-lowering term is multiplied by $t_{l}$ to get the scaled frequency as $\tilde{\omega}=\omega t_{l}$.
@@
Figure 1: Potential $V(x)$ with different asymmetry parameter value $c$ for a fixed barrier height of $h=1$.
@@
Figure 4: The initial probability distributions of the memory states for three values of $c$; $c=1$, the symmetric case, and $c=0.74$ and $0.56$, two asymmetric cases. (a) Probability distribution of the memory states after thermalization at $D=0.15$ corresponding to the initial equilibrium condition. The ratios of the memory states, 0 and $1$, are approximately as follows; $(50:50)$ for $c=1$, $(38:62)$ for $c=0.74$, and $(29:71)$ for $c=0.56$. (b) Initial non-thermalized $(50:50)$ distribution of the memory states for three corresponding values of $c$.
Figure 5: Stochastic trajectories of 100 particles (shown in orange, with a single trajectory highlighted in black) during the erasure process designed in bistable potentials with barrier height $h=1$, diffusion constant $D=0.1$, $\omega=0.01$, and barrier-lowering depth $Z=0.5$. Top row: Symmetric potential ($c=1$), showing (a) incomplete erasure at low tilt $A=0.2$, and (b) successful erasure at $A=1$ with equilibrium distribution of initial memory states. Middle row: Asymmetric potential ($c=0.2$), showing (c) incomplete erasure at $A=0.1$, and (d) successful erasure at $A=0.3$ with equilibrium distribution of initial memory states. Bottom row: Asymmetric potential ($c=0.2$), showing (e) incomplete erasure at $A=0.1$, and (f) successful erasure at $A=0.3$ with nonequilibrium distribution (50:50)(50{:}50) of initial memory states. For the cases with initial equilibrium states, the thermalization time is $t_{0}=314$. The ramp-up and ramp-down times for the tilting force are $\tau_{1}=298$ and $\tau_{2}=16$, respectively, for all cases.
@@ III Results and discussions
In the classical scenario, the Landauer bound is understood starting from an initial equal distribution to the memory states, 0 and 11, and then ultimate conversion to one unique memory state. In terms of the physical system, it implies initiating the process with equal occupancy of the Brownian particles in both wells, and then transferring all the particles through the erasure mechanism to a given well, which represents the final target memory state. Now, in the case of the symmetric bistable potential, this initial state is also the equilibrium state of the system before the start of the erasure protocol. Therefore, we can say that the Landauer limit is realized starting with the equilibrium memory. This initial equilibrium distribution can be achieved if we allow all Brownian trajectories to get adjusted with the thermal bath for a sufficiently long time [56, 15, 53, 57]. During this time, an individual Brownian particle gets settled in either of the two wells in the absence of any external force. When we consider an ensemble of such particles, the occupancy in the two wells will follow an equilibrium, more precisely, a steady-state distribution, in the long-time limit. Now, there is one important point to note. Although after this thermalization period, an equal equilibrium distribution in both the memory states is attained in the case of the symmetric potential, the thermally equilibrated distribution remains different for the asymmetric potentials. It also varies depending on the degree of asymmetry in the form of the potential. This is because after this equilibration time, the distribution follows from the steady state probability of occupancy of the two wells, which is now unequal for different wells due to their disparate widths in the asymmetric cases. Also, the thermalized distribution depends on the extent of asymmetry in the potential as the asymmetry parameter $c$ governs the relative widths of the two wells, thereby influencing the steady-state probability of their occupancy. Therefore, we can say that our choice of the potential of the asymmetric forms leads to the realization of the erasing process, starting from a condition when the two memory states occupy different phase space volumes, if we allow thermalization. However, the two memory states possess similar energy, initially, as the two minima of the potential correspond to the same energy value before the start of the erasure process. Consequently, we indicate that we start the erasing protocol with distributions of the memory in both the states, 0 and $1$, after the thermalization period. However, the corresponding occupancies of these two states are characteristics of the underlying potential. Therefore, one of the distinct features of our present study is that when we carry out the erasure process with a starting condition of equilibrium memory, the initial distribution of the memory states is unique based on the extent of asymmetry of the governing potential. We have presented representative initial equilibrium memory distributions for three conditions, one for the symmetric case $c=1$ and two asymmetric forms of the potential with $c=0.74$ and $0.56$ in Fig. 4(a). We perform our investigations on the erasure phenomenon with these initial conditions, for a range of asymmetric potential forms, to understand the effect of the asymmetry on it when we start with the equilibrium memory states.
@@
We infer that it is also important to analyze the effect of the asymmetry in results for the erasure mechanism when we start from the equally distributed initial memory states. This is because consideration of that initial state sets the limit of $k_{B}T\ln 2$ for the erasure of a classical bit of memory and interprets the Landauer bound. Therefore, we also carry out a separate set of investigations starting with the same condition of the equal initial distribution to the two memory states, irrespective of the structure of the potential. For the present case, we do not allow any thermalization time before the start of the protocol. We realize that this distribution will not conform to the thermalized distribution of the systems without any bias for the asymmetric cases, and can be regarded as the nonequilibrium initial condition of the memory states for those systems. Some representative probability distributions for the initial conditions with nonequilibrium memories have been illustrated in Fig.4(b). In this plot, the $c$ values for the potential have been chosen as the same as those considered in Fig.4(a) for the initial equilibrium condition scenarios. Performing the studies with these initial nonequilibrium distributions of memory states also provides us with the scope of comparisons of the findings between the cases of initial equilibrium and nonequilibrium memory conditions.